# FLARE: Towards Universal Dataset Purification against Backdoor Attacks

### v1

@@ Abstract
Deep neural networks (DNNs) are susceptible to backdoor attacks, where adversaries poison datasets with adversary-specified triggers to implant hidden backdoors, enabling malicious manipulation of model predictions. Dataset purification serves as a proactive defense by removing malicious training samples to prevent backdoor injection at its source. We first reveal that the current advanced purification methods rely on a latent assumption that the backdoor connections between triggers and target labels in backdoor attacks are simpler to learn than the benign features. We demonstrate that this assumption, however, does not always hold, especially in all-to-all (A2A) and untargeted (UT) attacks. As a result, purification methods that analyze the separation between the poisoned and benign samples in the input-output space or the final hidden layer space are less effective. We observe that this separability is not confined to a single layer but varies across different hidden layers. Motivated by this understanding, we propose FLARE, a universal purification method to counter various backdoor attacks. FLARE aggregates abnormal activations from all hidden layers to construct representations for clustering. To enhance separation, FLARE develops an adaptive subspace selection algorithm to isolate the optimal space for dividing an entire dataset into two clusters. FLARE assesses the stability of each cluster and identifies the cluster with higher stability as poisoned. Extensive evaluations on benchmark datasets demonstrate the effectiveness of FLARE against 22 representative backdoor attacks, including all-to-one (A2O), all-to-all (A2A), and untargeted (UT) attacks, and its robustness to adaptive attacks.
@@ I Introduction
Deep neural networks (DNNs) are widely deployed in mission-critical applications, including autonomous driving [18, 10], and face recognition [26, 27]. Currently, due to the complicated structure and large parameter scale of modern DNNs, training these models usually relies on large-scale datasets, typically from external sources (e.g., data markets and crowd-sourcing platforms).
However, recent studies reveal that using such external datasets introduces security risks, especially (poison-only) backdoor attacks [11, 5, 23, 47, 9]. Specifically, the backdoor adversaries poison a small subset of training data by embedding a predefined trigger (e.g., a subtle white square) and re-assigning their labels as the adversary-specified target label(s). As a result, all DNNs trained on these poisoned samples will learn a hidden backdoor, i.e., the latent and malicious connection between the triggers and the target labels. During the inference process of the backdoored model, the adversaries can activate its backdoor by implanting trigger patterns to maliciously change the predictions of any testing samples to the target labels. This attack is highly stealthy since the attacked model behaves normally on benign testing samples so that users cannot easily detect it simply based on the results of their local validation samples.
@@
Currently, researchers have investigated five representative defensive strategies against backdoor attacks. These strategies include: (1) dataset purification [39, 52, 15], (2) poison suppression [17, 8, 43], (3) model-level backdoor detection [45, 49, 46], (4) input-level backdoor detection [12, 37, 15], and (5) backdoor mitigation [28, 53, 50]. Among these strategies, poison suppression modifies the training process to reduce the impact of poisoned samples; model-level backdoor detection and mitigation identify or remove embedded backdoors in post-deployment; and input-level detection prevents backdoor activation by identifying malicious inputs during inference. In contrast, dataset purification acts as a proactive defense, identifying and removing poisoned samples from the dataset before training begins. This paper focuses on dataset purification, aiming to prevent backdoor creation at its source and precisely trace malicious samples to their origins.
In this paper, we first revisit existing advanced dataset purification methods. We reveal that their effectiveness relies on an implicit assumption: the connections between triggers and target labels in backdoor attacks are inherently simpler to learn than the benign features. This assumption enables these methods to achieve promising results by focusing primarily on input-output relationships. Specifically, (1) one line of work [20, 54] suggests that DNNs converge significantly faster on poisoned samples, indicating that the backdoor connections are easier to learn; (2) studies such as [6, 16] observe that poisoned samples often exhibit highly localized and small saliency regions. This suggests that backdoor triggers function as shortcut features, allowing the model to bypass learning complex, distributed features in benign data and thus simplifying the backdoor establishment; (3) research from [12, 15] leverages the scaled prediction consistency of poisoned samples, implying that DNNs overfit on distinctive, artificially crafted triggers, thereby simplifying backdoor formation. This assumption generally holds for typical all-to-one (A2O) backdoor attacks, as backdoor triggers are often simple and easy to learn. However, this assumption does not always hold for modern complicated backdoor attacks, especially for all-to-all (A2A) and untargeted (UT) backdoor attacks (as shown in Figure 1). As such, an intriguing question arises:
@@
The answer is in the positive! Inspired by the finding that DNN memorization is distributed across neurons in multiple hidden layers [32], we explore the distinctions between poisoned and benign samples throughout the model instead of simply the input-output relationships. Our analysis of A2A and UT backdoor attacks reveals a critical observation: poisoned and benign samples do not consistently separate within specific layers; instead, distinctions emerge across different hidden layers, varying with the attack types. Building on this insight, we propose a universal method for dataset purification, termed Full-spectrum Learning Analysis for Removing Embedded poisoned samples (FLARE). FLARE comprises two main stages: latent representation extraction and poisoned sample detection. In the first stage, FLARE constructs a comprehensive latent representation for each training sample by consolidating the abnormal values from all hidden layers’ feature maps. Specifically, FLARE first aligns all feature maps to a uniform scale (e.g., [0,1]) by leveraging the statistics of Batch Normalization (BN) layers. FLARE then extracts an abnormally large or small value from each feature map and consolidates these values across all hidden layers to construct the latent representation. In the second stage, FLARE detects poisoned samples through cluster analysis. In general, FLARE splits the entire dataset into two distinct clusters and identifies the cluster with higher cluster stability as poisoned. Specifically, FLARE first applies dimensionality reduction to reduce the computation consumption and improve clustering efficiency. FLARE then selects a stable subspace by adaptively excluding category-specific features from the last few hidden layers, isolating an optimal subspace where benign samples from various classes are still close together. FLARE evaluates cluster stability as the ‘density’ difference between the density level at which the cluster first appears and that where the cluster divides into smaller sub-clusters. FLARE finally identifies the cluster exhibiting higher stability as poisoned since poisoned samples tend to form compact clusters due to the sharing of the same trigger-related features.
@@ II-A Poison-only Backdoor Attacks
Targeted Attacks. Gu et al. [11] proposed the first poison-only backdoor attack, BadNets, where an adversary embeds an adversary-specified trigger into a few training samples and modifies their labels. This attack supports two primary modes: (1) all-to-one (A2O), where all poisoned samples are assigned a single target label, and (2) all-to-all (A2A), where poisoned samples from a class $i$ are relabeled as a different class label, typically the next consecutive class (i.e., $i+1$). Models trained on the poisoned dataset establish a backdoor connection between the trigger and target labels. Subsequent research aimed to make backdoor attacks more stealthy by developing invisible triggers [36] and employing label-consistent poisoning [44], where only samples from the target class are poisoned, thus avoiding label modification and bypassing manual inspection. Follow-up research further introduced more sophisticated trigger designs, including sample-specific triggers [25], sparse triggers [9], horizontal triggers [30], and asymmetric triggers[38], all specifically crafted to evade defenses.
Untargeted Attacks. Targeted attacks assign poisoned samples to adversary-specific labels, ensuring that the backdoored model consistently misclassifies the poisoned samples into particular classes. Recently, some pioneering work has discussed untargeted attacks [22, 51], where the goal is to make the predictions deviate from the true labels instead of approaching particular ones (i.e., target labels). For example, poisoned samples are reassigned with random incorrect labels [22], causing the backdoored model to misclassify the poisoned samples into incorrect classes. As a result, the predictions for all poisoned samples approximate a uniform distribution, making the attack more difficult to learn and detect.
@@ II-B Backdoor Defenses
Based on the stage at which they occur, existing defenses can be divided into five main categories: (1) dataset purification [39, 52, 15], which focuses on detecting and removing poisoned samples from a given suspicious dataset before model training, (2) poison suppression [17, 8, 43], which modifies the training process to limit the impact of poisoned samples, (3) model-level backdoor detection [45, 49, 46], which assesses whether a suspicious model contains hidden backdoors; (4) input-level backdoor detection [12, 37, 15], which identifies malicious inputs at inference; (5) backdoor mitigation [28, 53, 50], which directly removes backdoors after model development. This paper primarily focuses on dataset purification since it reduces backdoor threats from the source and can be easily used to mitigate model backdoors. We hereby provide an overview of the defences related to dataset purification and backdoor mitigation.
@@
Specifically, latent separability-based defenses exploited the detectable traces left by poisoned samples in the feature space. For example, Chen et al. [4] observed that, in the feature space of the final hidden layer, samples from the target class form two distinct clusters, with the smaller cluster identified as poisoned. Ma et al. [31] utilized high-order statistics (i.e., Gram matrix) to analyze the differences between poisoned and benign samples; Early convergence-based defenses relied on the observation that DNNs converge on poisoned samples more rapidly than on benign ones. During the early stages of training, the losses for poisoned samples quickly drop to near zero, while those for benign samples remain relatively high. For example, researchers in [20, 54] traped samples whose losses decreased more rapidly by using the local gradient ascent technique during the initial five training epochs. To mitigate class imbalance issues during this process, Gao et al. [8] refined the approach by selecting the lowest-loss samples within each class, rather than across the entire dataset; Dominant trigger-based defenses assume that backdoor triggers play a dominant role in DNN predictions. A backdoored model tends to learn an excessively strong signal for the backdoor trigger, such that even small, localized triggers can overpower other semantic features and dictate the model’s prediction. For example, Chou et al. [6] utilized model interpretability techniques (e.g., Grad-CAM [41]) to visualize salient regions of an input image, identifying highly localized and small regions as potential trigger areas. Similarly, Huang et al. [16] distilled minimal patterns from input images that influenced the model’s predictions and identified images with abnormally small patterns as poisoned; Perturbation consistency-based defenses assumed that poisoned samples are resistant to perturbations. For instance, Guo et al. [12] observed that poisoned samples exhibited prediction consistency under pixel-level amplification and proposed analyzing this consistency to distinguish poisoned samples. To address constraints in pixel values and insensitivity to amplification, Pal et al. [37] optimized a mask for selective pixel amplification; Qi et al. [39] analyzed the prediction consistency by unlearning the benign connections of the backdoored models; and Hou et al. [15] examined prediction consistency under unbounded weight-level amplification.
@@
However, in this paper, we find that all existing methods suffer from poor performance in some cases. Specifically, the first type of purification method generally only utilizes information at a specific layer (e.g., the last hidden layer) and can be easily bypassed by advanced attacks [38]. We will also show that the last three types of methods, however, relied on an underlying assumption that does not always hold, especially under A2A and UT attacks. How to design an effective dataset purification method is still an important open question.
@@
Backdoor Mitigation. This defense occurs in the post-development phase, aiming to remove implanted backdoors from attacked models. For instance, Li [28] pruned dormant neurons with benign inputs; Wu [48] pruned neurons that are sensitive to adversarial perturbations; Chai [3] applied weight-level pruning, which is more precise than neuron-level pruning, thereby preserving benign task performance; Li [21] adopted knowledge distillation to guide the fine-tuning of backdoored models. Instead of direct removal, researchers in [45, 53, 50] reverse-engineered suspected triggers and decouple them from target labels through unlearning and fine-tuning.
@@
Existing backdoor mitigation defenses have shown effectiveness against different attacks; however, Zhu et al. [55] demonstrated that the injected backdoors can persist and reactivate during inference, even after backdoor mitigation. This underscores the urgent need for a dataset purification method capable of defending against a broader range of attacks, including A2O, A2A, and UT attacks, to prevent backdoor creation at its source.
III Revisiting Existing Dataset Purification Methods
@@ III-A Preliminaries
The Main Pipeline of Poison-only Backdoor Attack. Let $\mathcal{D}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$ denotes a training set composed of $N$ i.i.d. samples. Each sample $({\bm{x}},y)$ is characterized by ${\bm{x}}\in\mathcal{X}=[0,1]^{d_{C}\times d_{W}\times d_{H}}$ and $y\in\mathcal{Y}=\{1,2,\ldots,K\}$. An adversary can generate a poisoned dataset $\hat{\mathcal{D}}$ by modifying a subset of benign samples (i.e., $\mathcal{D}_{s}$), i.e., $\hat{\mathcal{D}}=\mathcal{D}_{p}\cup\mathcal{D}_{b}$, where $\mathcal{D}_{b}=\mathcal{D}-\mathcal{D}_{s}\subset\mathcal{D}$, $\mathcal{D}_{p}=\{(\hat{{\bm{x}}},\hat{y})\mid\hat{{\bm{x}}}=\mathcal{G}_{X}({\bm{x}})),\hat{y}=\mathcal{G}_{Y}(y),({\bm{x}},y)\in\mathcal{D}_{s}\}$, and $\rho=|\mathcal{D}_{s}|/|\hat{\mathcal{D}}|$ is the poisoning rate. $\mathcal{G}_{X}:\mathcal{X}\rightarrow\mathcal{X}$, $\mathcal{G}_{Y}:\mathcal{Y}\rightarrow\mathcal{Y}$ are adversary-specified poisoned image generator and target label generator, respectively. For instance, in the BadNets attack [11], $\mathcal{G}_{X}(x)=(\bm{1}-{\bm{m}})\odot{\bm{x}}+{\bm{m}}\odot\bm{t}$, where ${\bm{m}}\in\{0,1\}^{d_{C}\times d_{W}\times d_{H}},\bm{t}\in\mathcal{X}$ is the malicious trigger. As for the $\mathcal{G}_{Y}(y)$, there are two primary attack paradigms: (1) targeted attacks and (2) untargeted attacks. Specifically, for the targeted attacks, poisoned images are re-labeled to the designated target labels: $\mathcal{G}_{Y}(y)=y_{t},y_{t}\in\mathcal{Y}$ in A2O attacks, and $\mathcal{G}_{Y}(y)=(y+1)\pmod{K}$ in A2A attacks. For the untargeted attacks, each poisoned image is re-labeled to a random label uniformly sampled from $\mathcal{Y}$ (i.e., $G_{Y}(y)\sim U(1,K)$).
@@
Figure 2: The benign accuracy (BA) and attack success rate (ASR) during the initial ten epochs of model training.
Threat Model of Dataset Purification. Similar to existing studies [4, 37], we assume that adversaries can freely poison the training dataset but can not manipulate the training process. Defenders aim to identify and filter out poisoned samples from a given third-party dataset. This is achieved through two main goals: effectiveness and generalizability. Effectiveness ensures that all poisoned samples are detected while minimizing the incorrect removal of benign samples. Generalizability guarantees that the defense is effective against various attack types, including A2O, A2A, and UT attacks. We assume that defenders have access to the poisoned training dataset and full control over the training process, but they have no knowledge of the backdoor attacks. In particular, our method requires less capacity for defenders compared to classical methods [39, 15] since it does not require additional benign local samples.
III-B Revisiting Strategy 1: Early Convergence
In general, the purification methods in [20, 17, 54] are designed based on the observation that DNNs converge faster on poisoned samples during the early stages of training, suggesting that the backdoor connections are easier to learn.
Settings. To validate the effectiveness of this purification strategy across different attack types, we conduct two representative backdoor attacks, i.e., BadNets [11] and WaNet [36], on the CIFAR-10 dataset using ResNet-18. They are the representatives of sample-agnostic and sample-specific attacks. Each attack includes three variants: A2O, A2A, and UT. For each attack, we calculate the benign accuracy (BA) and the attack success rate (ASR) during the initial ten epochs.

### v2

@@ Abstract
Deep neural networks (DNNs) are susceptible to backdoor attacks, where adversaries poison datasets with adversary-specified triggers to implant hidden backdoors, enabling malicious manipulation of model predictions. Dataset purification serves as a proactive defense by removing malicious training samples to prevent backdoor injection at its source. We first reveal that the current advanced purification methods rely on a latent assumption that the backdoor connections between triggers and target labels in backdoor attacks are simpler to learn than the benign features. We demonstrate that this assumption, however, does not always hold, especially in all-to-all (A2A) and untargeted (UT) attacks. As a result, purification methods that analyze the separation between the poisoned and benign samples in the input-output space or the final hidden layer space are less effective. We observe that this separability is not confined to a single layer but varies across different hidden layers. Motivated by this understanding, we propose FLARE, a universal purification method to counter various backdoor attacks. FLARE aggregates abnormal activations from all hidden layers to construct representations for clustering. To enhance separation, FLARE develops an adaptive subspace selection algorithm to isolate the optimal space for dividing an entire dataset into two clusters. FLARE assesses the stability of each cluster and identifies the cluster with higher stability as poisoned. Extensive evaluations on benchmark datasets demonstrate the effectiveness of FLARE against 22 representative backdoor attacks, including all-to-one (A2O), all-to-all (A2A), and untargeted (UT) attacks, and its robustness to adaptive attacks. Codes are available at BackdoorBox and backdoor-toolbox.
@@ I Introduction
Deep neural networks (DNNs) are widely deployed in mission-critical applications, including autonomous driving kong2020physgan ; grigorescu2020survey , and face recognition li2016mutual ; li2014common . Currently, due to the complicated structure and large parameter scale of modern DNNs, training these models usually relies on large-scale datasets, typically from external sources (e.g., data markets).
However, recent studies reveal that using such external datasets introduces security risks li2023detecting ; zhang2024badrobot ; 10795257 ; zhang2025improving ; song2025pb . In particular, backdoor attacks have been identified as a significant threat gu2017badnets ; li2022backdoor ; wei2024pointncbw ; gao2024backdoor ; luo2025just . Specifically, the backdoor adversaries poison a small subset of training data by embedding a predefined trigger (e.g., a subtle white square) and re-assigning their labels as the adversary-specified target label(s). As a result, all DNNs trained on these poisoned samples will learn a hidden backdoor, i.e., the latent and malicious connection between the triggers and the target labels. During the inference process of the backdoored model, the adversaries can activate its backdoor by implanting trigger patterns to maliciously change the predictions of any samples to the target labels. This attack is highly stealthy since the attacked model behaves normally on benign samples so that users cannot easily detect it simply based on the results of their local validation samples.
@@
Currently, researchers have investigated five representative defensive strategies against backdoor attacks. These strategies include: (1) dataset purification qi2023towards ; yao2024reverse ; hou2024ibd , (2) poison suppression huang2022backdoor ; gao2023backdoor ; tang2023setting , (3) model-level backdoor detection wang2019neural ; xiang2023umd ; wang2024mm , (4) input-level backdoor detection guo2023scale ; pal2024backdoor ; hou2024ibd , and (5) backdoor mitigation liu2018fine ; zeng2022adversarial ; xu2024towards . Among these strategies, poison suppression modifies the training process to reduce the impact of poisoned samples; model-level backdoor detection and mitigation identify or remove embedded backdoors in post-development; and input-level detection prevents backdoor activation by identifying malicious inputs during inference. In contrast, dataset purification acts as a proactive defense, identifying and removing poisoned samples from the dataset before training begins. This paper focuses on dataset purification, aiming to prevent backdoor creation at its source and precisely trace malicious samples to their origins.
In this paper, we first revisit existing advanced dataset purification methods. We reveal that their effectiveness relies on an implicit assumption: the connections between triggers and target labels in backdoor attacks are inherently simpler to learn than the benign features. This assumption enables these methods to achieve promising results by focusing primarily on input-output relationships. Specifically, (1) one line of work li2021anti ; zhang2023backdoor suggests that DNNs converge significantly faster on poisoned samples, indicating that the backdoor connections are easier to learn; (2) studies such as chou2020sentinet ; huang2023distilling observe that poisoned samples often exhibit highly localized and small saliency regions. This suggests that backdoor triggers function as shortcut features, allowing the model to bypass learning complex, distributed features in benign data and thus simplifying the backdoor establishment; (3) research from guo2023scale ; hou2024ibd leverages the scaled prediction consistency of poisoned samples, implying that DNNs overfit on distinctive, artificially crafted triggers, thereby simplifying backdoor formation. This assumption generally holds for typical all-to-one (A2O) backdoor attacks, as backdoor triggers are often simple and easy to learn. However, this assumption does not always hold for modern complicated backdoor attacks, especially for all-to-all (A2A) and untargeted (UT) backdoor attacks (as shown in Figure 1). As such, an intriguing question arises:
@@
The answer is in the positive! Inspired by the finding that DNN memorization is distributed across neurons in multiple hidden layers maini2023can , we explore the distinctions between poisoned and benign samples throughout the model instead of simply the input-output relationships. Our analysis of A2A and UT backdoor attacks reveals a critical observation: poisoned and benign samples do not consistently separate within specific layers; instead, distinctions emerge across different hidden layers, varying with the attack types. Building on this insight, we propose a universal method for dataset purification, termed Full-spectrum Learning Analysis for Removing Embedded poisoned samples (FLARE). FLARE comprises two main stages: latent representation extraction and poisoned sample detection. In the first stage, FLARE constructs a comprehensive latent representation for each training sample by consolidating the abnormal values from all hidden layers’ feature maps. Specifically, FLARE first aligns all feature maps to a uniform scale (e.g., [0,1]) by leveraging the statistics of Batch Normalization (BN) layers. FLARE then extracts an abnormally large or small value from each feature map and consolidates these values across all hidden layers to construct the latent representation. In the second stage, FLARE detects poisoned samples through cluster analysis. In general, FLARE splits the entire dataset into two distinct clusters and identifies the cluster with higher cluster stability as poisoned. Specifically, FLARE first applies dimensionality reduction to reduce the computation consumption and improve clustering efficiency. FLARE then selects a stable subspace by adaptively excluding category-specific features from the last few hidden layers, isolating an optimal subspace where benign samples from various classes are still close together. FLARE evaluates cluster stability as the ‘density’ difference between the density level at which the cluster first appears and that where the cluster divides into smaller sub-clusters. FLARE finally identifies the cluster exhibiting higher stability as poisoned since poisoned samples tend to form compact clusters due to the sharing of the same trigger-related features.
@@ II Related Work
DNNs are inherently vulnerable to various security threats due to their high model complexity, large parameter space, and tendency to overfit. These vulnerabilities make DNNs susceptible to a range of attacks, including adversarial attacks gao2025pixels ; zhang2024does , membership inference he2025labelonly ; li2024yes , and backdoor attacks gu2017badnets ; gao2020backdoor . Among them, backdoor attacks represent an emerging but highly threatening security vulnerability gu2017badnets ; li2022backdoor . In such attacks, an adversary implants a malicious connection, known as a backdoor, between a predefined trigger and an adversary-specified target label. This backdoor remains dormant for benign inputs and only activates when the trigger appears, making backdoor attacks stealthy.
@@ II-A Poison-only Backdoor Attacks
Targeted Attacks. Gu et al. gu2017badnets proposed the first poison-only backdoor attack, BadNets, where an adversary embeds an adversary-specified trigger into a few training samples and modifies their labels. This attack supports two primary modes: (1) all-to-one (A2O), where all poisoned samples are assigned a single target label, and (2) all-to-all (A2A), where poisoned samples from a class $i$ are relabeled as a different class label, typically the next consecutive class (i.e., $i+1$). Models trained on the poisoned dataset establish a backdoor connection between the trigger and target labels. Subsequent research aimed to make backdoor attacks more stealthy by developing invisible triggers nguyen2021wanet and employing label-consistent poisoning turner2019label , where only samples from the target class are poisoned, thus avoiding label modification and bypassing manual inspection. Follow-up research further introduced more sophisticated triggers, including sample-specific triggers li2021invisible , sparse triggers gao2024backdoor , horizontal triggers ma2023horizontal , and asymmetric triggers qi2023revisiting , crafted to evade defenses.
Untargeted Attacks. Targeted attacks assign poisoned samples to adversary-specific labels, ensuring that the backdoored model consistently misclassifies the poisoned samples into particular classes. Recently, some pioneering work has discussed untargeted attacks li2022untargeted ; xue2023untargeted , where the goal is to make the predictions deviate from the true labels instead of approaching particular ones (i.e., target labels). For example, poisoned samples are reassigned with random incorrect labels li2022untargeted , causing the backdoored model to misclassify the poisoned samples into incorrect classes. As a result, the predictions for all poisoned samples approximate a uniform distribution, making the attack more difficult to learn and detect.
@@ II-B Backdoor Defenses
Based on the stage at which they occur, existing defenses can be divided into five main categories: (1) dataset purification qi2023towards ; yao2024reverse ; gao2025try , which focuses on detecting and removing poisoned samples from a given suspicious dataset before model training, (2) poison suppression huang2022backdoor ; gao2023backdoor ; tang2023setting , which modifies the training process to limit the impact of poisoned samples, (3) model-level backdoor detection wang2019neural ; xiang2023umd ; wang2024mm , which assesses whether a suspicious model contains hidden backdoors; (4) input-level backdoor detection gao2019strip ; li2023ntd ; luo2025just , which identifies malicious inputs at inference; (5) backdoor mitigation zeng2022adversarial ; li2024purifying ; xu2024towards , which directly removes backdoors after model development. This paper primarily focuses on dataset purification since it reduces backdoor threats from the source and can be easily used to mitigate model backdoors. We hereby provide an overview of the defences related to dataset purification and backdoor mitigation.
@@
Specifically, latent separability-based defenses exploited the detectable traces left by poisoned samples in the feature space. For example, Chen et al. chen2018detecting observed that, in the feature space of the final hidden layer, samples from the target class form two distinct clusters, with the smaller cluster identified as poisoned. Ma et al. ma2022beatrix utilized high-order statistics (i.e., Gram matrix) to analyze the differences between poisoned and benign samples; Early convergence-based defenses relied on the observation that DNNs converge on poisoned samples more rapidly than on benign ones. During the early stages of training, the losses for poisoned samples quickly drop to near zero, while those for benign samples remain relatively high. For example, researchers in li2021anti ; zhang2023backdoor traped samples whose losses decreased more rapidly by using the local gradient ascent technique during the initial five training epochs; Dominant trigger-based defenses assume that backdoor triggers play a dominant role in DNN predictions. For example, Chou et al. chou2020sentinet utilized model interpretability techniques (e.g., Grad-CAM selvaraju2017grad ) to visualize salient regions of an input image, identifying highly localized and small regions as potential trigger areas. Similarly, Huang et al. huang2023distilling distilled minimal patterns from input images that influenced the model’s predictions and identified images with abnormally small patterns as poisoned; Perturbation consistency-based defenses assumed that poisoned samples are resistant to perturbations. For instance, Guo et al. guo2023scale observed that poisoned samples exhibited prediction consistency under pixel-level amplification and proposed analyzing this consistency to distinguish poisoned samples. To address constraints in pixel values and insensitivity to amplification, Pal et al. pal2024backdoor optimized a mask for selective pixel amplification; Qi et al. qi2023towards analyzed the prediction consistency by unlearning the benign connections of the backdoored models; and Hou et al. hou2024ibd examined prediction consistency under unbounded weights.
@@
However, in this paper, we find that all existing methods suffer from poor performance in some cases. Specifically, the first type of purification method generally only utilizes information at a specific layer (e.g., the last hidden layer) and can be easily bypassed by advanced attacks qi2023revisiting . We will also show that the last three types of methods, however, relied on an underlying assumption that does not always hold, especially under A2A and UT attacks. Most recently, a concurrent work (i.e., Telltale gao2025try ), offered a novel and effective solution for detecting poisoned samples in the training process. It gathered loss trajectories of samples after model convergence and applied clustering analysis to identify poisoned clusters. It is applicable to non-classification tasks, mitigating the reliance on clean data and addressing inconsistencies in defense effectiveness. However, it focuses mainly on the A2O and clean-label attack modes, overlooking other attack modes such as A2A and UT. How to design an effective and comprehensive dataset purification method is still an important open question.
@@
Backdoor Mitigation. This defense occurs in the post-development phase, aiming to remove implanted backdoors from attacked models. For instance, Li liu2018fine pruned dormant neurons with benign inputs; Wu wu2021adversarial pruned neurons that are sensitive to adversarial perturbations; Chai chai2022one applied weight-level pruning, which is more precise than neuron-level pruning, thereby preserving benign task performance; Li li2021neural adopted knowledge distillation to guide the fine-tuning of backdoored models. Instead of direct removal, researchers in wang2019neural ; zeng2022adversarial ; xu2024towards reverse-engineered triggers and decouple them from target labels through unlearning and fine-tuning.
@@
Existing backdoor mitigation defenses have shown effectiveness against different attacks; however, Zhu et al. zhu2024breaking demonstrated that the injected backdoors can persist and reactivate during inference, even after backdoor mitigation. This underscores the urgent need for a dataset purification method capable of defending against a broader range of attacks, including A2O, A2A, and UT attacks, to prevent backdoor creation at its source.
III Revisiting Dataset Purification
@@ III-A Preliminaries
The Main Pipeline of Poison-only Backdoor Attack. Let $\mathcal{D}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$ denotes a training set composed of $N$ i.i.d. samples. Each sample $({\bm{x}},y)$ is characterized by ${\bm{x}}\in\mathcal{X}=[0,1]^{d_{C}\times d_{W}\times d_{H}}$ and $y\in\mathcal{Y}=\{1,2,\ldots,K\}$. An adversary can generate a poisoned dataset $\hat{\mathcal{D}}$ by modifying a subset of benign samples (i.e., $\mathcal{D}_{s}$), i.e., $\hat{\mathcal{D}}=\mathcal{D}_{p}\cup\mathcal{D}_{b}$, where $\mathcal{D}_{b}=\mathcal{D}-\mathcal{D}_{s}\subset\mathcal{D}$, $\mathcal{D}_{p}=\{(\hat{{\bm{x}}},\hat{y})\mid\hat{{\bm{x}}}=\mathcal{G}_{X}({\bm{x}})),\hat{y}=\mathcal{G}_{Y}(y),({\bm{x}},y)\in\mathcal{D}_{s}\}$, and $\rho=|\mathcal{D}_{s}|/|\hat{\mathcal{D}}|$ is the poisoning rate. $\mathcal{G}_{X}:\mathcal{X}\rightarrow\mathcal{X}$, $\mathcal{G}_{Y}:\mathcal{Y}\rightarrow\mathcal{Y}$ are adversary-specified poisoned image generator and target label generator, respectively. For instance, in the BadNets attack gu2017badnets , $\mathcal{G}_{X}(x)=(\bm{1}-{\bm{m}})\odot{\bm{x}}+{\bm{m}}\odot\bm{t}$, where ${\bm{m}}\in\{0,1\}^{d_{C}\times d_{W}\times d_{H}},\bm{t}\in\mathcal{X}$ is the malicious trigger. As for the $\mathcal{G}_{Y}(y)$, there are two primary attack paradigms: (1) targeted attacks and (2) untargeted attacks. Specifically, for the targeted attacks, poisoned images are re-labeled to the designated target labels: $\mathcal{G}_{Y}(y)=y_{t},y_{t}\in\mathcal{Y}$ in A2O attacks, and $\mathcal{G}_{Y}(y)=(y+1)\pmod{K}$ in A2A attacks. For the untargeted attacks, each poisoned image is re-labeled to a random label uniformly sampled from $\mathcal{Y}$ (i.e., $G_{Y}(y)\sim U(1,K)$).
@@
Figure 2: The benign accuracy (BA) and attack success rate (ASR) during the initial ten epochs of model training. These curves show the model’s convergence behavior, with rapid increases in ASR indicating that the corresponding backdoor connection is easier to learn, while slower rises suggest more challenging ones, especially under A2A and UT attacks.
Threat Model of Dataset Purification. Similar to existing studies chen2018detecting ; pal2024backdoor , we assume that adversaries can freely poison the training dataset but can not manipulate the training process. Defenders aim to identify and filter out poisoned samples from a given third-party dataset. This is achieved through two main goals: effectiveness and generalizability. Effectiveness ensures that all poisoned samples are detected while minimizing the incorrect removal of benign samples. Generalizability guarantees that the defense is effective against various attack types, including A2O, A2A, and UT attacks. We assume that defenders have access to the poisoned training dataset and full control over the training process, but they have no knowledge of the backdoor attacks. In particular, our method requires less capacity for defenders compared to classical methods qi2023towards ; hou2024ibd since it does not require additional benign local samples.
III-B Revisiting Strategy 1 (Early Convergence)
In general, the purification methods in li2021anti ; huang2022backdoor ; zhang2023backdoor are designed based on the observation that DNNs converge faster on poisoned samples during the early stages of training, suggesting that the backdoor connections are easier to learn.
Settings. To validate the effectiveness of this purification strategy across different attack types, we conduct two representative backdoor attacks, i.e., BadNets gu2017badnets and WaNet nguyen2021wanet , on the CIFAR-10 dataset using ResNet-18. They are the representatives of sample-agnostic and sample-specific attacks. Each attack includes three variants: A2O, A2A, and UT. For each attack, we calculate the benign accuracy (BA) and the attack success rate (ASR) during the initial ten epochs.